a carnot engine is operated between two reservoirs of temperature 1770c and 770c if the engine receives 4200 J of heat energy from the source is each cycle, calculate the amount of heat rejected to the sink in each cycle. calculate the efficiency of the engine.

Grade:11

2 Answers

Arun
6 years ago
Dear student

efficiency of carnot engine = 1 – T2/ T1
 = 1 – (1043/ 2043)
 = 1 – 0.510
 = 0.49
hence efficiency = 49%
efficiency  = w/Q
i hope now you can calculate heat rejected.
In case of any query, please feel free to ask.
 
Regards
Arun (askIITians forum expert)
Nani yeggireddy
6 years ago
T1 =177+273k=450k T2= 77 +273k =350kQ1=4200J,Q2=?Formula Q2/Q1=T2/T1Q2=350×4200/350=3266.67JEffeciency of engine =1-T2/T1=1-350/450=0.2222=22.22%